# Deep clustering using adversarial net based clustering loss

### v1

@@ I Introduction
Deep neural network such as the autoencoder is applied to many signal processing domains such as speech and audio processing [1, 2], image clustering [3, 4] and medical data processing [5, 6]. While the latent space of autoencoder is well suited for dimension reduction through reconstruction loss [7], the latter is not optimized for clustering/ classification since class labels cannot be used in the reconstruction loss [8]. A recent autoencoder technique known as the Euclidean distance based clustering (ABC) [7], utilize class label information in neural network by introducing a loss function known as the deep clustering loss. The goal is to minimizes the Euclidean distance in the latent space between samples and partitioning learnt by clustering algorithms e.g. Kmeans or Gaussian mixture model (GMM). In other words, the latent space of an autoencoder learnt using reconstruction loss will look different when we further perform K-means on the latent space. A recent probabilistic approach to ABC uses KL divergence (KLD) [4] and assumes that both the variational autoencoder (VAE) [9] latent space and the unsupervised clustering approach are Gaussian distributed. We can compare the similarities and differences between deep clustering and VAE to better understand the former. The neural network of VAE and deep clustering are optimized by backpropagating samples in the latent space to update the encoder weights. Also, both deep clustering and VAE share a similar goal of modeling each input image as a sample draw of the Gaussian distribution in the latent space. However, their Gaussian distributions are different. VAE enforce samples in the latent space to be closely distributed to a Gaussian distribution with continuous mean $\mu$ and variance $\sigma$ i.e. $z\sim\mu+\sigma\cdot\mathcal{N}(0,1)$. Whereas the latent space of deep clustering approaches a Gaussian mixture model (GMM) and each Gaussian has a set of fixed parameters i.e. $z\mathcal{\sim N}(\eta_{k},\tau_{k})$, $\forall k\in K$. The goal of VAE is to enforce all class samples in the latent space to be distributed and centered around a zero mean and unit variance Gaussian, $\mathcal{N}(0,1)$. The goal of deep clustering is to enforce class samples in the latent space to be closely distributed to their respective nearest Gaussian clusters i.e. $N(\eta^{*},\tau^{*})$ and nearest cluster is denoted by $\eta^{*}=\arg\max_{k}\eta_{k}$ and vice versa for $\tau^{*}$. Thus, only deep clustering is associated with clustering, making it more suitable for unsupervised classification.
@@ II Background
II-A Deep clustering using Euclidean distance loss
The simplest form of deep clustering uses the Euclidean distance loss in the Autoencoder based clustering (ABC) [7, 4] approach. In ABC, the deep clustering objective optimizes the neural network to be affected by Kmeans clustering or $L_{2}$ loss function in eqn (1). On the contrary, traditional reconstruction loss or $L_{1}$ does not utilize any information from Kmeans.
$\begin{array}[]{c}\underset{w,b}{min}\;\left\{Reconstruction+ABC\right\}\\ \\ =\underset{w,b}{min}\;\left\{L_{1}+\lambda\cdot L_{2}\right\}\\ \\ =\underset{w,b}{min}\;\left\|T-y\right\|^{2}-\lambda\left\|h_{z}-\eta^{*}\right\|^{2}\\ \\ =\underset{w,b}{max}\;-\frac{1}{2}\left(T-y\right)^{2}-\lambda\left\{-\frac{1}{2}\left(h_{z}-\eta^{*}\right)^{2}\right\}\end{array}$ (1)
From the perspective of weight update, the effect of cluster mean (i.e. $\eta^{*}$) on the neural network can be seen initially in the gradient, $\frac{\delta L_{2}}{\delta h_{z}}=\lambda\left(h_{z}-\eta^{*}\right)$. Assuming an autoencoder architecture below that uses a single hidden layer neural network for the encoder,
$\begin{array}[]{c}h_{j}=f_{j}\left(\sum_{i}w_{ij}\cdot x_{i}\right)\\ h_{z}=f_{z}\left(\sum_{j}w_{jz}\cdot h_{j}\right)\\ h_{k}=f_{k}\left(\sum_{z}w_{zk}\cdot h_{z}\right)\\ y_{l}=f_{l}\left(\sum_{k}w_{kl}\cdot h_{k}\right)\\ \\ \end{array}$ (2)
the effect or presence of Kmeans i.e. $\frac{\delta L_{2}}{\delta h_{z}}$ can be found in the weight update in eqn (3) when performing backpropagation
$\begin{array}[]{c}\Delta w_{j\mu}=\frac{\delta L1}{\delta w_{jz}}+\lambda\left(\frac{\delta L_{2}}{\delta h_{z}}\cdot\frac{\delta h_{z}}{\delta w_{jz}}\right)\\ \Delta w_{ij}=\frac{\delta L1}{\delta w_{ij}}+\lambda\left(\frac{\delta L_{2}}{\delta h_{z}}\cdot\frac{\delta h_{z}}{\delta h_{j}}\cdot\frac{\delta h_{j}}{\delta w_{ij}}\right)\\ \\ \end{array}$ (3)
II-B Probabilistic variant of deep clustering
@@
A) From ABC to KLD to JSD:
We show that under certain condition (i.e. unit variance assumption), ABC is identical to KLD and thus related to JSD. Both KLD and JSD are also under the family of the $f$-divergence.
B) Adversarial net as JSD:
We approximate JSD using adversarial net. This is possible since adversarial net approaches JSD when the training becomes optimal.
Our main goal is to find a relationship between adversarial net and deep clustering. First, we must first identify the relationship between KLD and ABC as seen in [4]. KLD measures the probabilistic distance between two distributions. The distributions of the latent space and deep clustering space [^1] We refer to deep clustering space as $p(z\mid\theta)$ and the latent space as $q(z\mid x)$. We refer to deep clustering space as the latent space partitioned using GMM. are defined as $q(z\mid x)$ and $p(z\mid\theta)$ respectively. The latent space is defined using the re-parameterization trick in [9] i.e. $z=\mu+\varrho\cdot\epsilon$. The deep clustering space is defined as $p(z\mid\theta)$, whereby $z=\left\{z^{(n)}\right\}_{n=1}^{N}\in\mathbb{R}^{Z}$ and $\theta=\{\eta,\tau,\zeta\}$, which are the mean $\eta$, precision $\tau$ and assignment parameters $\boldsymbol{\zeta}$ of GMM for $K$ number of clusters and $N$ number of samples. Also, $\eta=\left\{\eta_{k}\right\}_{k=1}^{K}\in\mathbb{R}^{Z},\tau=\left\{\tau_{k}\right\}_{k=1}^{K}\in\mathbb{R}^{Z},\zeta_{k}^{(n)}\in\{0,1\}$ and $\zeta^{(n)}$ is a $1-of-K$ vector. Thus, the KLD based clustering loss in [4] as follows:
$\begin{array}[]{c}D_{KL}\left(\;p(z\mid\theta)\parallel q(z\mid x)\;\right)\\ \\ p(z\mid\theta)=\prod_{k=1}^{K}\mathcal{N}(z\mid\eta_{k},\left(\tau_{k}\right)^{-1})^{\zeta_{k}}\\ q(z\mid x)=\mu+\sigma\cdot\mathcal{N}(0,1)\end{array}$ (4)
The problem is how do we define KLD in terms of two Gaussians, instead of a Gaussian and a GMM in eqn (4), which is intractable. To overcome this, we re-express the GMM term in $p(z\mid\theta)$ as
@@
We define $p(z\mid\theta^{*})$ as the $k^{th}$ Gaussian component of Kmeans or GMM that generates sample $z$ in the latent space. i.e. we use $\zeta$ to compute $\eta^{*}$ and $\tau^{*}$ (refer to GMM in Appendix). Thus, the KLD of GMM and latent space in eqn (4) can now become a KLD between two Gaussian distributions. Also, a closed formed equation is available for the latter.
@@
The relationship between KLD and ABC can therefore be explained: If we discard the second order terms i.e. $\varrho^{2}$ and $\tau$ in eqn (6), the KLD reverts back to the original ABC loss [7, 14] in eqn (1). We refer to Lemma 1. for further details.
$D_{KL}\left(\mathcal{N}(z\mid\eta^{*})\parallel\mathcal{N}(z\mid\mu)\right)=\frac{1}{2}\left(\eta^{*}-\mu\right)^{2}$ (7)
Lemma 1.
Relating ABC [7] to KLD: We show that under assumption of “unit variance”, the $KL$D between the encoder latent space and GMM reverts back to the ABC loss.
Proof:
For a $D_{KL}$ between two Gaussian distributions, there is a unique closed-form expression available. When we assume unit variance i.e. $\tau=\sigma=1$, the $D_{KL}$ reverts back to the original Euclidean distance based clustering loss in eqn (1) as follows ∎
$\begin{array}[]{c}D_{KL}\left(p(z\mid\theta)\parallel q(z\mid x)\right)\;\;s.t.\>\>\{\tau=\sigma=1\}\\ \\ =D_{KL}\left(\mathcal{N}(z_{n}\mid\eta^{*},\left(\tau^{*}\right){}^{-1})\parallel\mathcal{N}(z_{n}\mid\mu,\sigma)\right)\\ \\ =\ln\tau^{*}+\ln\sigma+\frac{\left(\tau^{*}\right)^{-1}+\left(\eta^{*}-\mu\right)^{2}}{2\sigma^{2}}-\frac{1}{2}\\ \\ =\frac{1}{2}\left(\eta^{*}-\mu\right)^{2}\end{array}$ (8)
To relate the KLD to JSD, we simply recall JSD as the averaging between two KLDs in eqn (9). We next discuss how to relate adversarial net to JSD.
$\begin{array}[]{c}D_{JS}\left(\>p(z\mid\theta^{*})\parallel q(z\mid x)\>\right)\\ \\ =\frac{1}{2}D_{KL}\left(p\parallel\frac{p+q}{2}\right)+\frac{1}{2}D_{KL}\left(q\parallel\frac{p+q}{2}\right)\end{array}$ (9)
II-C Adversarial net approaches JSD
@@
The problem with using JSD for deep clustering is that there is rarely a closed form solution available for a “JSD between two Gaussian distributions”. GAN overcome this by employing an adversarial training procedure that approximates $D_{JS}\left(p_{data}\parallel p_{g}\right)$ at optimality, where $p_{data}$ is the real data distribution and $p_{g}$ is the generated data distribution as follows [10]:
$\begin{array}[]{c}D_{JS}\left(p_{data}\parallel p_{g}\right)\\ \\ =\underset{x\sim p_{data}}{E}\left[\log D_{G}(x)\right]+\underset{z\sim p(z)}{E}\left[\log\left(1-D_{G}\left\{G(z)\right\}\right)\right]\\ \\ =\underset{x\sim p_{data}}{E}\left[\log D_{G}(x)\right]+\underset{x\sim p_{g}}{E}\left[\log\left(1-D_{G}\left\{x\right\}\right)\right]\\ \\ \end{array}$ (10)
Unlike the discriminator of GAN in the original space, the discriminator of DCAN works in the latent space. Despite that, we can easily reformulate GAN into DCAN as follows:
@@
$\begin{array}[]{c}D_{JS}\left(p(z\mid\theta^{*})\parallel q(z\mid x)\right)\\ \\ =\underset{z\sim p(z\mid\theta^{*})}{E}\left[\ln D(z)\right]+\underset{x\sim p_{data}}{E}\left[\ln\left(1-D\left\{G(x)\right\}\right)\right]\\ \\ =\underset{z\sim p(z\mid\theta^{*})}{E}\left[\ln D(z)\right]+\underset{z\sim q(z\mid x)}{E}\left[\ln\left(1-D\left\{z\right\}\right)\right]\end{array}$ (11)
@@
Whereby $E_{z\sim p(z\mid\theta^{*})}\left[\cdot\right]$ refers to the expectation function where the sample is distributed from deep clustering space and $E_{z\sim q(z\mid x)}\left[\cdot\right]$ refers to the expectation function where the sample is distributed from the latent space. We refer to Lemma 2 for the claim on eqn (11).
Lemma 2.
The adversarial net based loss of DCAN can be shown to approach JSD at optimum.
Proof:
For the sake of brevity we refer to $p(z\mid\theta^{*})$ and $q(z\mid x)$ as $p$ and $q$ respectively. The optimal discriminator occurs when $G$ is fixed, i.e. $D=\frac{p}{p+q}$ Substituting $D$, we define the LHS and RHS below. Lastly, if we subject $p=q=1$ on both sides, thus we can validate the claim on eqn (11).
$\begin{array}[]{c}\boldsymbol{LHS}:E_{z\sim p}\left[\ln D(z)\right]+E_{x\sim p_{data}}\left[\ln\left(1-D\left\{G(x)\right\}\right)\right]\cdots\\ \;\;s.t.\>\>\{D=\frac{p}{p+q}\}\\ \\ =E_{z\sim p}\left[\log D(z)\right]+E_{z\sim q}\left[\log\left(1-D(z)\right)\right]\;\;s.t.\>\>\{D=\frac{p}{p+q}\}\\ \\ =E_{z\sim p}\left[\ln\frac{p}{p+q}\right]+E_{z\sim q}\left[\ln\frac{q}{p+q}\right]\\ \\ =\int_{z}\>p\ln\frac{p}{p+q}+q\ln\frac{q}{p+q}\>dz\\ \\ \boldsymbol{RHS}:D_{JS}\left(p\parallel q\right)=\frac{1}{2}\int\>p\ln\frac{2\cdot p}{p+q}+q\ln\frac{2\cdot p}{p+q}\>dz\\ \\ =\frac{1}{2}\int_{z}\>p\left\{\ln\frac{p}{p+q}+\ln 2\right\}+q\left\{\ln\frac{q}{p+q}+\ln 2\right\}\>dz\\ \\ \end{array}$ (12)
Thus, $LHS\leq 2RHS-2\log 2$. ∎
III Training the Proposed Method, DCAN
@@
We can visualize $D_{JS}\left(\>p(z\mid\theta^{*})\parallel q(z\mid x)\>\right)$ using Fig. 1. The discriminator tries to distinguish samples from deep clustering space, $z\sim p(z\mid\theta^{*})$ versus samples from the encoder output, $z\sim q(z\mid x)$. We train DCAN by using the two backward paths shown by the colored dashed lines in Fig 1. We attach a pseudocode of DCAN in Algorithm 1.
@@ III-A Training the discriminator
$\begin{array}[]{c}D=y^{T}(1-y)^{1-T}\\ \\ L_{1}=\ln D(z)\\ \\ L_{2}=\ln\left[1-D(G(x))\right]\end{array}$ (13)
@@
Neural Network R: We perform forward pass using eqn (14) and stochastic gradient ascent (SGA) updates using eqn (15). As we seek sampling from $p(z\mid\theta^{*})$ instead of $q(z\mid x)$, we set $T=0$ for $D$. From eqn (13), the gradient of $L_{2}$ i.e. $\delta L_{2}/\delta y$ is shown in eqn (16).
$\begin{array}[]{cc}h_{j}=\sigma_{j}\left(\sum_{i=1}^{I}w_{ij}\cdot x_{i}\right),&h_{f}=\sigma_{f}\left(\sum_{\varsigma=1}^{Z}w_{\varsigma f}\cdot z_{\varsigma}\right)\\ \\ z_{\varsigma}=\sigma_{\varsigma}\left(\sum_{j=1}^{J}w_{j\varsigma}\cdot h_{j}\right),&y_{d}=\sigma_{d}\left(\sum_{f=1}^{F}w_{fd}\cdot h_{f}\right)\end{array}$ (14)
$\begin{array}[]{cc}w_{fd}=w_{fd}+\gamma\triangle w_{fd},&\triangle w_{fd}=\frac{\delta L_{2}}{\delta y}\times\frac{\delta y}{\delta\sigma_{d}}\times\frac{\delta\sigma_{d}}{\delta w_{fd}}\\ \\ w_{\varsigma f}=w_{\varsigma f}+\gamma\triangle w_{\varsigma f},&\triangle w_{\varsigma f}=\frac{\delta L_{2}}{\delta y}\times\frac{\delta y}{\delta\sigma_{d}}\times\frac{\delta\sigma_{d}}{\delta h_{f}}\cdots\\ &\times\frac{\delta h_{f}}{\delta\sigma_{f}}\times\frac{\delta\sigma_{f}}{\delta w_{\varsigma f}}\end{array}$ (15)
$\begin{array}[]{c}\frac{\delta L_{2}}{\delta y}=\frac{\delta}{\delta y}\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{x^{\left(n\right)}\sim p_{data}}{E}\left[\ln\left(1-D\left(G(x^{\left(n\right)})\right)\right)\right]\right\}\end{array}$ (16)
@@
$\begin{array}[]{c}\eta_{k}=\frac{1}{N}\sum_{n=1}^{N}\;z^{(n)}\in k\\ \\ \tau_{k}=\left(\frac{1}{N}\sum_{n=1}^{N}\;\left(z^{(n)}-\eta_{k}\right)^{2}\right)^{-1}\in k\\ \\ \zeta^{(n)}=\arg\max_{k}\left\{\;\ln p(z^{(n)}\mid\eta_{k},\tau_{k})\;\right\}\forall k\end{array}$ (17)
The samples in the latent space displaces each time the encoder weight changes. Thus, GMM has to be re-computed each time the weight is updated. After sampling from GMM i.e. $z\sim p(z\mid\theta^{*})$ and setting $T=1$ to $D$, we perform forward pass using eqn (18) and SGA using eqn (19). Using $L_{1}$ in eqn (13), $\delta L_{1}/\delta y$ is computed in eqn (20).
$\begin{array}[]{c}h_{f}=\sigma_{f}\left(\sum_{\varsigma=1}^{Z}w_{\varsigma f}\cdot\left\{z\sim p(z\mid\theta^{*})\right\}\right)\\ \\ y_{d}=\sigma_{d}\left(\sum_{f=1}^{F}w_{fd}\cdot h_{f}\right)\end{array}$ (18)
$\begin{array}[]{cc}w_{fd}=w_{fd}+\gamma\triangle w_{fd},&\triangle w_{fd}=\frac{\delta L_{1}}{\delta w_{y}}\times\frac{\delta y}{\delta\sigma_{d}}\times\frac{\delta\sigma_{d}}{\delta w_{fd}}\\ \\ w_{\varsigma f}=w_{\varsigma f}+\gamma\triangle w_{\varsigma f},&\triangle w_{\varsigma f}=\frac{\delta L_{1}}{\delta y}\times\frac{\delta y}{\delta\sigma_{d}}\times\frac{\delta\sigma_{d}}{\delta h_{f}}\cdots\\ &\times\frac{\delta h_{f}}{\delta\sigma_{f}}\times\frac{\delta\sigma_{f}}{\delta w_{\varsigma f}}\end{array}$ (19)
$\begin{array}[]{c}\frac{\delta L_{1}}{\delta y}=\frac{\delta}{\delta y}\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{z^{\left(n\right)}\sim p(z|\theta^{*})}{E}\left[\ln D\left(z^{\left(n\right)}\right)\right]\right\}\end{array}$ (20)
III-B Training the encoder

### v2

@@ I Introduction
Deep neural network such as the autoencoder is applied to many signal processing domains such as speech and audio processing [1, 2], image clustering [3, 4] and medical data processing [5, 6]. While the latent space of autoencoder is well suited for dimension reduction through reconstruction loss [7], the latter is not optimized for clustering/ classification since class labels cannot be used in the reconstruction loss [8]. A recent autoencoder technique known as the Euclidean distance based clustering (ABC) [7], utilize class label information in neural network by introducing a loss function known as the deep clustering loss. The goal is to minimizes the Euclidean distance in the latent space between samples and partitioning learnt by clustering algorithms e.g. Kmeans or Gaussian mixture model (GMM). In other words, the latent space of an autoencoder learnt using reconstruction loss will look different when we further perform K-means on the latent space. A recent probabilistic approach to ABC uses KL divergence (KLD) [4, 9] and assumes that both the variational autoencoder (VAE) [10] latent space and the unsupervised clustering approach are Gaussian distributed. We can compare the similarities and differences between deep clustering and VAE to better understand the former. The neural network of VAE and deep clustering are optimized by backpropagating samples in the latent space to update the encoder weights. Also, both deep clustering and VAE share a similar goal of modeling each input image as a sample draw of the Gaussian distribution in the latent space. However, their Gaussian distributions are different. VAE enforce samples in the latent space to be closely distributed to a Gaussian distribution with continuous mean $\mu$ and variance $\sigma$ i.e. $z\sim\mu+\sigma\cdot\mathcal{N}(0,1)$. Whereas the latent space of deep clustering approaches a Gaussian mixture model (GMM) and each Gaussian has a set of fixed parameters i.e. $z\mathcal{\sim N}(\eta_{k},\tau_{k})$, $\forall k\in K$. The goal of VAE is to enforce all class samples in the latent space to be distributed and centered around a zero mean and unit variance Gaussian, $\mathcal{N}(0,1)$. The goal of deep clustering is to enforce class samples in the latent space to be closely distributed to their respective nearest Gaussian clusters i.e. $N(\eta^{*},\tau^{*})$ and nearest cluster is denoted by $\eta^{*}=\arg\max_{k}\eta_{k}$ and vice versa for $\tau^{*}$. Thus, only deep clustering is associated with clustering, making it more suitable for unsupervised classification.
@@ II Background
II-A Kmeans based deep clustering
The simplest form of deep clustering uses the Euclidean distance loss in the Autoencoder based clustering (ABC) [7] approach. In ABC, the deep clustering objective optimizes the neural network to be affected by Kmeans clustering or $\mathcal{L}_{1}$ loss function in eqn (1). On the contrary, traditional reconstruction loss or $\mathcal{L}_{2}$ does not utilize any information from Kmeans.
$\begin{array}[]{c}\underset{w,b}{min}\;\left\{\mathcal{L}_{1}+\lambda\cdot\mathcal{L}_{2}\right\}\\ =\underset{w,b}{max}\;-\frac{1}{2}\left(T-y\right)^{2}-\lambda\left\{-\frac{1}{2}\left(h_{z}-\eta^{*}\right)^{2}\right\}\end{array}$ (1)
II-B KLD based deep clustering
@@
A) From ABC to KLD to JSD: We show that under certain condition (i.e. unit variance assumption), ABC is identical to KLD and thus related to JSD. Both KLD and JSD are also under the family of the $f$-divergence.
B) Adversarial net as JSD: We approximate JSD using adversarial net. This is possible since adversarial net approaches JSD when the training becomes optimal [11].
Our main goal is to find a relationship between adversarial net and deep clustering. First, we must first identify the relationship between KLD and ABC as seen in [4]. KLD measures the probabilistic distance between two distributions. The distributions of the latent space and deep clustering space [^1] We refer to deep clustering space as $p(z\mid\theta)$ and the latent space as $q(z\mid x)$. We refer to deep clustering space as the latent space partitioned using GMM. are defined as $q(z\mid x)$ and $p(z\mid\theta)$ respectively. The latent space is defined using the re-parameterization trick in [10] i.e. $z=\mu+\varrho\cdot\epsilon$. The deep clustering space is defined as $p(z\mid\theta)$, whereby $z=\left\{z^{(n)}\right\}_{n=1}^{N}\in\mathbb{R}^{Z}$ and $\theta=\{\eta,\tau,\zeta\}$, which are the mean $\eta$, precision $\tau$ and assignment parameters $\boldsymbol{\zeta}$ of GMM for $K$ number of clusters and $N$ number of samples. Also, $\eta=\left\{\eta_{k}\right\}_{k=1}^{K}\in\mathbb{R}^{Z},\tau=\left\{\tau_{k}\right\}_{k=1}^{K}\in\mathbb{R}^{Z},\zeta_{k}^{(n)}\in\{0,1\}$ and $\zeta^{(n)}$ is a $1-of-K$ vector. Thus, KLD based clustering loss in [4] as follows:
$\begin{array}[]{c}D_{KL}\left(\;p(z\mid\theta)\parallel q(z\mid x)\;\right),\\ p(z\mid\theta)=\prod_{k=1}^{K}\mathcal{N}(z\mid\eta_{k},\left(\tau_{k}\right)^{-1})^{\zeta_{k}},\\ q(z\mid x)=\mu+\sigma\cdot\mathcal{N}(0,1)\end{array}$ (2)
The problem is how do we define KLD in terms of two Gaussians, instead of a Gaussian and a GMM in eqn (4), which is intractable. To overcome this, we re-express the GMM term in $p(z\mid\theta)$ as follows
@@
We define $p(z\mid\theta^{*})$ as the $k^{th}$ Gaussian component of Kmeans or GMM that generates sample $z$ in the latent space. i.e. we use $\zeta$ to compute $\eta^{*}$ and $\tau^{*}$ (refer to GMM in Appendix). Thus, the KLD of GMM and latent space in eqn (2) can now become a KLD between two Gaussian distributions in eqn (4). More importantly, a closed formed equation is available for the latter.
@@
The relationship between KLD and ABC can further be explained: If we discard the second order terms i.e. $\varrho^{2}$ and $\tau$ in eqn (4), the KLD loss reverts back to the original ABC loss [7, 14] in eqn (1).
III Proposed JSD based deep clustering
@@
The problem with using JSD for deep clustering is that there is rarely a closed form solution available for a “JSD between two Gaussian distributions”. GAN [11] overcome this by employing an adversarial training procedure that approximates $D_{JS}\left(p_{data}\parallel p_{g}\right)$ at optimality, where $p_{data}$ is the real data distribution and $p_{g}$ is the generated data distribution. Unlike the discriminator of GAN in the original space, the discriminator of DCAN works in the latent space. Despite that, we can reformulate GAN into DCAN as follows:
@@
$\begin{array}[]{c}D_{JS}\left(p(z\mid\theta^{*})\parallel q(z\mid x)\right)\\ =\underset{z\sim p(z\mid\theta^{*})}{E}\left[\ln D(z)\right]+\underset{x\sim p_{data}}{E}\left[\ln\left(1-D\left\{G(x)\right\}\right)\right]\\ =\underset{z\sim p(z\mid\theta^{*})}{E}\left[\ln D(z)\right]+\underset{z\sim q(z\mid x)}{E}\left[\ln\left(1-D\left\{z\right\}\right)\right]\end{array}$ (5)
@@
Whereby $E_{z\sim p(z\mid\theta^{*})}\left[\cdot\right]$ refers to the expectation function where the sample is distributed from deep clustering space and $E_{z\sim q(z\mid x)}\left[\cdot\right]$ refers to the expectation function where the sample is distributed from the latent space.
IV Training the Proposed Method
@@
We can visualize $D_{JS}\left(\>p(z\mid\theta^{*})\parallel q(z\mid x)\>\right)$ using Fig. 1. The discriminator tries to distinguish samples from deep clustering space, $z\sim p(z\mid\theta^{*})$ versus samples generated from the encoder output, $z\sim q(z\mid x)$. We train DCAN by using the two backward paths shown by the colored dashed lines in Fig 1. We attach a pseudocode of DCAN in Algorithm 1.
@@ IV-A Training the discriminator
$\begin{array}[]{c}D=y^{T}(1-y)^{1-T}\\ L_{1}=\ln D(z)\\ L_{2}=\ln\left[1-D(G(x))\right]\end{array}$ (6)
@@
Neural Network R: We perform weight update using eqn (7) via stochastic gradient ascent (SGA). As we seek sampling from $p(z\mid\theta^{*})$ instead of $q(z\mid x)$, we set $T=0$ for $D$.
$\begin{array}[]{c}L_{2}=\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{x^{\left(n\right)}\sim p_{data}}{E}\left[\ln\left(1-D\left(G(x^{\left(n\right)})\right)\right)\right]\right\}\end{array}$ (7)
@@
$\begin{array}[]{c}\eta_{k}=\frac{1}{N}\sum_{n=1}^{N}\;z^{(n)}\in k\\ \tau_{k}=\left(\frac{1}{N}\sum_{n=1}^{N}\;\left(z^{(n)}-\eta_{k}\right)^{2}\right)^{-1}\in k\\ \zeta^{(n)}=\arg\max_{k}\left\{\;\ln p(z^{(n)}\mid\eta_{k},\tau_{k})\;\right\}\forall k\end{array}$ (8)
Since the samples in the latent space displaces each time the encoder weight changes, GMM has to be re-computed each time the weight is updated. After sampling from GMM i.e. $z\sim p(z\mid\theta^{*})$ and setting $T=1$ to $D$, we perform weight update using eqn (9).
$\begin{array}[]{c}L_{1}=\left\{\frac{1}{N}\sum_{n=1}^{N}\;\underset{z^{\left(n\right)}\sim p(z|\theta^{*})}{E}\left[\ln D\left(z^{\left(n\right)}\right)\right]\right\}\end{array}$ (9)
IV-B Training the encoder